OpenStudy (anonymous):

How do I solve and check for the linear equation 4x+1=9

OpenStudy (anonymous):

4x + 1= 9 4x = 9 - 1 4x = 8 x = 8/4 To check, just plug in the value u found for x and see if the left side matches the right side.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

np